The question is: find three consecutive intengers whose sum is 11 less than 5 times the smallest of the three.
I think the equation would be x+y+z=5x-11 but how do I solve it?
Help!

Answer by scott8148(6628)

you are on the right track

consecutive integers means y is one more than x and z is one more than y ... so x+(x+1)+(x+2)=5x-11

3x+3=5x-11 ... 14=2x ... 7=x